Please help me solve this equation: ((four to the power of sixteen divided by four to the power of negative 2 equals))

Answer by stanbon(75887)

((four to the power of sixteen divided by four to the power of negative 2 equals))
----
Rule: When dividing, subtract the exponents.
--------------------------
4^16 / 4^(-2) = 4^(16--2) = 4*18
=========================================
Cheers,
Stan H.